OpenStudy (anonymous):

what is the simplified trig expression for csc theta sin theta-sin^2 theta?

OpenStudy (raden):

we knowed that csc(theta) = 1/sin(theta) so, csc theta sin theta-sin^2 theta = 1/sin(theta) * sin(theta) -sin^2 theta = 1 - sin^2 theta

OpenStudy (raden):

then use the identity sin^2(theta) + cos^2(theta) = 1 we get 1 - sin^2 theta = cos^2(theta)

OpenStudy (anonymous):

I have that so far but wasn't sure what to do with 1-sin^2theta

OpenStudy (anonymous):

first of all you would have to know that \[\csc \theta = \frac{ 1 }{ \sin \theta } \]